# Technical insights and legal considerations for advancing federated learning in bioinformatics

### v1

@@ 2.6 Privacy-enhancing techniques
Homomorphic encryption (HE; Gentry, 2009) is a cryptographic technique that enables computations to be performed directly on encrypted data (ciphertexts) without requiring decryption. The outcome of operations on ciphertexts matches the result of performing the same operations on the corresponding non-encrypted values (plaintexts) when decrypted. HE can be either fully homomorphic (FHE), which allows for arbitrary computations, or partially homomorphic (PHE), which supports only a specific subset of mathematical operations. For instance, the Paillier PHE scheme (Paillier, 1999) only supports additive operations on encrypted data. FHE requires considerable computational resources for encryption and decryption. PHE is less flexible but computationally more efficient, making it a common choice in practical applications.
@@
Secure multiparty computation (SMPC; Zhao et al., 2019) is a peer-to-peer protocol allowing multiple parties to compute a function over their data collaboratively, similarly to Figure 2 (centre). Each data holder divides their data into random shares and distributes them among all parties in the consortium, thus ensuring that no single party can access the complete data set. The shares are then combined during the computation process, often with the assistance of a server, to produce the correct result while preserving data privacy. SMPC ensures high security with exact results and keeps data private throughout the computation process. However, SMPC is computationally intensive and requires peer-to-peer communication, leading to high communication overhead. Its complexity also increases with the number of participants, limiting scalability.
@@
Another approach to securing FL is using an aggregator and a compensator server in a centralised two-server topology (Figure 2, right; Nasirigerdeh et al., 2021). Each client adds a noise pattern to their local data, sharing the former with the compensator (which aggregates all noise patterns) and the latter with the aggregator (which aggregates the noisy data and trains the model). The aggregator then obtains the overall noise pattern from the compensator and removes it from the aggregated noisy data, allowing for denoised model training. This two-server approach is efficient: it requires neither extensive computation in the clients nor peer-to-peer communication. However, it makes infrastructure more complex and requires trust in both servers not to collude to compromise the privacy of individual contributions.
@@
Unlike the above methods, which are encryption-based methods ensuring data confidentiality during transmission or storage, differential privacy, another popular technique for data-protection in federated learning, is not an encryption system but rather a technique that focuses on privacy by ensuring that the output of data analysis does not leak sensitive information about the underlying dataset. Differential privacy (DP; Ficek et al., 2021) achieves this by a mathematical framework designed to enable analyses to remain statistically consistent regardless of whether any specific individual’s data is included or excluded. This property guarantees that sensitive information about individuals cannot be inferred from the results up to a preset “privacy budget” worth of operations. DP is typically implemented by introducing noise into the data (Schein et al., 2019; Cai et al., 2021), weight clipping in the training process (Abadi et al., 2016; Jayaraman and Evans, 2019) or predictions (Nissim et al., 2007; Dwork and Feldman, 2018) to obfuscate individual contributions. The amount of noise must be carefully calibrated to balance predictive accuracy and privacy within the analysis: too little noise undermines privacy, and too much reduces performance. This effect is more pronounced within specific subgroups underrepresented in the training set (Bagdasaryan et al., 2019).
@@ 3 Federated learning in bioinformatics
Most FL literature focuses on general algorithms and is motivated by applications other than bioinformatics, such as digital twins for smart cities (Ramu et al., 2022), smart industry (Zhang et al., 2021) and open banking and finance (Long et al., 2020). Even the clinical literature mainly focuses on different types of data and issues (Dayan et al., 2021; van Rooden et al., 2024). Here, we highlight and discuss notable examples of FL designed specifically for bioinformatics. They are all in the early stages of development, so their reliability, reproducibility, and scalability are open questions. However, they hint at the potential of FL to perform better than meta-analysis and single-client analyses on real-world data, comparing favourably to centralised data analyses in which data are pooled in a central location while addressing data sharing and use concerns (Zolotareva et al., 2021; Xu et al., 2020).
@@ 3.1 Proteomics and differential gene expression
Proteomics studies the complex protein dynamics that govern cellular processes and their interplay with physiological and pathological states, such as cancer (Maes et al., 2015), to improve risk assessment, treatment selection and patient monitoring. Differential expression analyses focus specifically on comparing expression levels across different conditions, tissues, or cell types to identify genes with statistically significant differences (Rodriguez-Esteban and Jiang, 2017).
In addition to the issues discussed in Section 2, FL in proteomics must overcome the challenge of integrating data from different platforms (Rieke et al., 2020) while accounting for imbalanced samples and batch effects. Cai et al. (2022) produced a federated implementation of DEqMS (FedProt; Zhu et al., 2020) for variance estimation in mass spectrometry-based data that successfully identifies top differentially-abundant proteins in two real-world data sets using label-free quantification and tandem mass tags.
Zolotareva et al. (2021) implemented a federated limma voom pipeline (Law et al., 2014) on top of HyFed (Nasirigerdeh et al., 2021), which uses the aggregator-compensator two-server topology we described earlier. This approach was showcased on two extensive RNA-seq data sets, proving robust to heterogeneity across clients and batch effects. Hannemann et al. (2024) trained a federated deep-learning model for cell type classification using both Flower and TFF and different architectures, with similar results.
@@ 3.2 Genome-wide association studies
Genome-wide association studies (GWAS) aim to identify genomic variants statistically associated with a qualitative (say, a case-control label) or quantitative trait (say, body mass index). These studies mainly use regression models, which can be largely trained using general-purpose federated regression implementations with minor modifications to address scalability and correct for population structure (see, for instance, Kolobkov et al., 2024).
@@
As an alternative, Cho et al. (2024) build on REGENIE (Mbatchou et al., 2021) to avoid using a genomic relatedness matrix and increase the scalability of GWAS while using MPC and HE to secure the data. Despite the overhead introduced by the encryption, this approach is efficient enough to work on a cohort of 401k individuals from the UK Biobank and 90 million single-nucleotide polymorphisms (SNPs) in less than 5 hours.
@@ 3.3 Single-cell RNA sequencing
Single-cell RNA sequencing (scRNA-seq) measures gene expression at the cellular level rather than aggregating it at the tissue level as in bulk RNA sequencing, identifying the distinct expression profiles of individual cell populations within tissues (Hwang et al., 2018; Papalexi and Satija, 2018).
Wang et al. (2024) developed scFed, a unified FL framework integrating four algorithms for cell type classification from scRNA-seq data: the ACTINN neural network (Ma and Pellegrini, 2020), explicitly designed for this task; a linear support vector machine; XGBoost based on Li et al. (2023b); and the GeneFormer transformer (Theodoris et al., 2023). They evaluated scFed on eight data sets evenly distributed among 2–20 clients, suggesting that the federated approach has a predictive accuracy comparable to that obtained by pooling the data and better than that in individual clients. However, the overhead during training increases with the number of clients, limiting the scalability to larger consortia.
@@ 3.4 Multi-omics
Proteomics, genomics, and transcriptomics capture different aspects of biological processes. Integrating large data sets from different omics offers deeper insights into their underlying mechanisms (Civelek and Lusis, 2014). Vertical FL allows multiple parties to combine various features of the same patients into multimodal omics data sets without exposing sensitive information (Liu et al., 2024). For instance, Wang et al. (2023) trained a deep neural network with an adaptive optimisation module for cancer prognosis evaluation from multi-omics data. The neural network performs feature selection while the adaptive optimisation module prevents overfitting, a common issue in small high-dimensional samples (Rajput et al., 2023). This method performs better than a single-omic analysis, but the improvement in predictive accuracy is strongly model-dependent. Another example is Danek et al. (2024), who built a diagnostic model for Parkinson’s disease: they provided a reproducible setup for evaluating several multi-omics models trained on pre-processed, harmonised and artificially horizontally federated data using Flower. Their study identifies a general but not substantial reduction in FL performance compared to centrally trained models, which increases with the number of clients and is variably affected by client heterogeneity.
@@ 3.5 Medical imaging
Medical imaging studies the human body’s interior to diagnose abnormalities in its anatomy and physiology from digital images such as those obtained by radiography, magnetic resonance and ultrasound devices (Suetens, 2017). It is the most common application of FL in the medical literature (Chowdhury et al., 2022). As a result, protocols for image segmentation and diagnostic prediction are well documented. Notable case studies target breast cancer (Roth et al., 2020), melanomas (Haggenmüller et al., 2024), cardiovascular disease (Linardos et al., 2022), COVID-19 (Yang et al., 2021; Dayan et al., 2021).
Machine learning applications that use medical imaging data typically face challenges such as incomplete or inaccurate labelling and the normalisation of images from different scanners and different protocols. Bdair et al. (2022) explored a federated labelling scheme in which clients produced ground-truth labels for skin lesions in a privacy-preserving manner, improving classification accuracy. Yan et al. (2023) also proposed an efficient scheme to use data sets mainly comprising unlabelled images, focusing on chest X-rays. Furthermore, Jiang et al. (2022) apply FL to learn a harmonised feature set from heterogeneous medical images, improving both the classification and segmentation of histology and MRI scans.
@@ 3.6 Ready-to-use FL tools for bioinformatics
The need for user-friendly FL implementations of common bioinformatics workflows has driven the creation of secure collaborative analysis tools (Berger and Cho, 2019; Froelicher et al., 2021; Wan et al., 2022). Two notable examples are sfkit and FeatureCloud.
@@
The sfkit framework (Mendelsohn et al., 2023) facilitates federated genomic analyses by implementing GWAS, principal component analysis (PCA), genetic relatedness and a modular architecture to complement them as needed. It provides a web interface featuring a project bulletin board, chat functions, study parameter configurations and results sharing. State-of-the-art cryptographic tools for privacy preservation based on SMPC and HE ensure data protection (Mouchet et al., 2020).
@@
FeatureCloud (Matschinske et al., 2023) is an integrated solution from which end users without programming experience can build custom workflows. It provides modules to run on the clients and servers in the consortium. Unlike sfkit, FeatureCloud allows users publish applications in its app store, including regression models, random forests and neural networks. Developers must also document how privacy guarantees are implemented in their apps.
4 How to conduct federation of specific operations in bioinformatics
@@
This section provides practical insights to help a reader interested in building a federated and secure analogue of an existing bioinformatics algorithm. We focus on horizontal FL with the centralised topology from Figure 2 (left). Consider different clients, each possessing a local data set , where . Each data set contains samples, denoted as , where represents the sample index, and represents the features for each sample. We denote a row (column) of the matrix as (). This describes a distributed data set of observations:
@@
The following sections assume that an FL consortium has been established, the necessary infrastructure is operational, and an appropriate FL framework has been selected and installed. It is also assumed that a secure aggregation protocol has been chosen, such as those described in Section 2.6 and Figure 4. The choice of a specific secure aggregation protocol may depend on several factors, including technology and infrastructure (e.g., the availability of a specific FL topology that drives the choice), privacy risks, or scalability concerns, as discussed in Section 2.6. In the following sections, we provide a general overview of sum-based mathematical operations built upon a secure aggregation protocol, as well as operations involving federated averaging (FedAvg; McMahan et al., 2017).
@@
Coding examples using Flower (Beutel et al., 2020) are available in our GitHub repository (https://github.com/IDSIA/FL-Bioinformatics). We chose Flower because it has a shallow learning curve for new FL users and provides a good balance between simplicity and flexibility when implementing custom algorithms. Riedel et al. (2024) also identified Flower as a promising framework because it has a large, active, and growing community of developers and scientists and because of the extensive tutorials and documentation. In our examples, secure summation is performed using the secure aggregation protocol SecAgg+ (Bell et al., 2020). This protocol combines encryption with SMPC, using a multiparty approach in which each client interacts with only a subset of the others. It is particularly suitable for several FL contexts, as it is robust to client dropout and highly scalable. In particular, a relevant aspect of the bioinformatics domain is that it scales linearly with the size of the vectors to be aggregated (Li et al., 2021).
@@ 4.1 Sum-based computations
Let be real numbers stored by individual clients. We define the secure sum of these numbers, performed through the selected secure aggregation protocol, as . We can build on this simple secure sum to construct a wide range of operations. However, note that as the complexity of operations increases, the amount of information revealed to the server may also increase. Sum-based operations include:

### v4

@@ 3.6 Privacy-enhancing techniques
Homomorphic encryption (HE; 48) is a cryptographic technique that enables computations to be performed directly on encrypted data (ciphertexts) without requiring decryption. The outcome of operations on ciphertexts matches the result of performing the same operations on the corresponding non-encrypted values (plaintexts) when decrypted. HE can be either fully homomorphic (FHE), which allows for arbitrary computations, or partially homomorphic (PHE), which supports only a specific subset of mathematical operations. For instance, the Paillier PHE scheme (49) only supports additive operations on encrypted data. FHE requires considerable computational resources for encryption and decryption. PHE is less flexible but computationally more efficient, making it a common choice in practical applications.
@@
Secure multiparty computation (SMPC; 50) is a peer-to-peer protocol allowing multiple parties to compute a function over their data collaboratively, similarly to Figure 2 (centre). Each data holder divides their data into random shares and distributes them among all parties in the consortium, thus ensuring that no single party can access the complete data set. The shares are then combined during the computation process, often with the assistance of a server, to produce the correct result while preserving data privacy. SMPC ensures high security with exact results and keeps data private throughout the computation process. However, SMPC is computationally intensive and requires peer-to-peer communication, leading to high communication overhead. Its complexity also increases with the number of participants, limiting scalability.
@@
Another approach to securing FL is using an aggregator and a compensator server in a centralised two-server topology (Figure 2, right; 25). Each client adds a noise pattern to their local data, sharing the former with the compensator (which aggregates all noise patterns) and the latter with the aggregator (which aggregates the noisy data and trains the model). The aggregator then obtains the overall noise pattern from the compensator and removes it from the aggregated noisy data, allowing for denoised model training. This two-server approach is efficient: it requires neither extensive computation in the clients nor peer-to-peer communication. However, it makes infrastructure more complex and requires trust in both servers not to collude to compromise the privacy of individual contributions.
@@
Unlike the above methods, which are encryption-based methods ensuring data confidentiality during transmission or storage, differential privacy, another popular technique for data-protection in federated learning, is not an encryption system but rather a technique that focuses on privacy by ensuring that the output of data analysis does not leak sensitive information about the underlying dataset. Differential privacy (DP; 51) achieves this through a mathematical framework designed to ensure analyses remain statistically consistent, regardless of whether any specific individual’s data is included or excluded. This property guarantees that sensitive information about individuals cannot be inferred from the results up to a preset ”privacy budget” worth of operations. DP is typically implemented by introducing noise into the data (52, 53), weight clipping in the training process (54, 55) or predictions (56, 57) to obfuscate individual contributions. The amount of noise must be carefully calibrated to balance predictive accuracy and privacy within the analysis: too little noise undermines privacy, and too much reduces performance. This effect is more pronounced within specific subgroups underrepresented in the training set (58).
@@ 4 Federated learning in bioinformatics
Most FL literature focuses on general algorithms and is motivated by applications other than bioinformatics, such as digital twins for smart cities (59), smart industry (60) and open banking and finance (61). Even the clinical literature mainly focuses on different types of data and issues (11, 62). Here, we highlight and discuss notable examples of FL designed specifically for bioinformatics, summarised in Table 2. They are all in the early stages of development, so their reliability, reproducibility, and scalability are open questions. However, they hint at the potential of FL to perform better than meta-analysis and single-client analyses on real-world data, comparing favourably to centralised data analyses where data are pooled in a central location while addressing data sharing and use concerns (20, 15).
@@ 4.1 Proteomics and differential gene expression
Proteomics studies the complex protein dynamics that govern cellular processes and their interplay with physiological and pathological states, such as cancer (63), to improve risk assessment, treatment selection and patient monitoring. Differential expression analyses focus specifically on comparing expression levels across different conditions, tissues, or cell types to identify genes with statistically significant differences (64).
In addition to the issues discussed in Section 3, FL in proteomics must overcome the challenge of integrating data from different platforms (65) while accounting for imbalanced samples and batch effects. Cai et al. (66) produced a federated implementation of DEqMS (FedProt; 67) for variance estimation in mass spectrometry-based data that successfully identifies top differentially-abundant proteins in two real-world data sets using label-free quantification and tandem mass tags.
Zolotareva et al. (20) implemented a federated limma voom pipeline (68) on top of HyFed (25), which uses the aggregator-compensator two-server topology we described earlier. This approach was demonstrated on two extensive RNA-seq data sets, proving robust to heterogeneity across clients and batch effects. Hannemann et al. (69) trained a federated deep-learning model for cell type classification using both Flower and TFF and different architectures, with similar results.
@@ 4.2 Genome-wide association studies
Genome-wide association studies (GWAS) aim to identify genomic variants statistically associated with a qualitative (say, a case-control label) or quantitative trait (say, body mass index). These studies mainly use regression models, which can be largely trained using general-purpose federated regression implementations with minor modifications to address scalability and correct for population structure (see, for instance, 70).
@@
As an alternative, Cho et al. (74) built on REGENIE (75) to avoid using a genomic relatedness matrix and increase the scalability of GWAS while using MPC and HE to secure the data. Despite the overhead introduced by the encryption, this approach is efficient enough to work on a cohort of 401k individuals from the UK Biobank and 90 million single-nucleotide polymorphisms (SNPs) in less than 5 hours.
| Field | Application | FL Methods | Data | References |
| Proteomics and Differential Gene Expression | Variance estimation, gene expression, cell type classification | FedDEqMS, FedProt, HyFed with limma voom, DL with Flower and TFF | Mass spectrometry, RNA-seq, 1-10k individuals and 10-100M biomarkers | Cai et al. (66) Zolotareva et al. (20) Hannemann et al. (69) Nasirigerdeh et al. (25) |
| GWAS | Association testing, scalable regression | FedGLMM, federated GRM estimator, FedGMMAT, REGENIE with MPC/HE | SNPs, genotype and phenotype data, 2,5–275k individuals and 0.5–38M SNPs | Li et al. (71) Wang et al. (72) Li et al. (73) Cho et al. (74) |
| Single-cell RNA-seq | Cell type classification | scFed: ACTINN, SVM, XGBoost, GeneFormer | scRNA-seq from 2-55k cells and 1-2k genes | Wang et al. (76) Li et al. (77) |
| Multi-omics | Prognosis (cancer), diagnostics (Parkinson’s) | Vertical FL, adaptive neural networks, benchmarking with Flower | Genomics, transcriptomics, proteomics, 100–1200 individuals and 100-700 features | Wang et al. (78) Danek et al. (79) |
| Medical Imaging | Classification, segmentation, semi-supervised training | Federated labelling, harmonised feature learning | MRI, X-rays, histology images, 5–71k scans | Bdair et al. (80) Yan et al. (81) Jiang et al. (82) Haggenmüller et al. (83) Linardos et al. (84) Yang et al. (85) |
| Specialised Tools | FL software for bioinformatics workflows | sfkit, FeatureCloud | All the data above | Mendelsohn et al. (19) Matschinske et al. (86) Berger and Cho (87) Froelicher et al. (88) |
@@ 4.3 Single-cell RNA sequencing
Single-cell RNA sequencing (scRNA-seq) measures gene expression at the cellular level, rather than aggregating it at the tissue level as in bulk RNA sequencing, and identifies the distinct expression profiles of individual cell populations within tissues (89, 90).
Wang et al. (76) developed scFed, a unified FL framework integrating four algorithms for cell type classification from scRNA-seq data: the ACTINN neural network (91), explicitly designed for this task; a linear support vector machine; XGBoost based on Li et al. (77); and the GeneFormer transformer (92). They evaluated scFed on eight data sets evenly distributed among 2–20 clients, suggesting that the federated approach has a predictive accuracy comparable to that obtained by pooling the data and better than that in individual clients. However, the overhead during training increases with the number of clients, limiting the scalability to larger consortia. More recently, Bakhtiari et al. (93) introduced FedscGen, a federated implementation of scGen (94), a variational autoencoder-based method for batch effect correction. FedscGen employs secure SMPC for privacy-preserving aggregation and achieves results that closely match those obtained under centralised training.
@@ 4.4 Multi-omics
Proteomics, genomics, and transcriptomics capture different aspects of biological processes. Integrating large data sets from different omics offers deeper insights into their underlying mechanisms (95). Vertical FL allows multiple parties to combine various features of the same patients into multimodal omics data sets without exposing sensitive information (96). For instance, Wang et al. (78) trained a deep neural network with an adaptive optimisation module for cancer prognosis evaluation from multi-omics data. The neural network performs feature selection while the adaptive optimisation module prevents overfitting, a common issue in small high-dimensional samples (97). This method performs better than a single-omic analysis, but the improvement in predictive accuracy is strongly model-dependent. Another example is Danek et al. (79), who built a diagnostic model for Parkinson’s disease: they provided a reproducible setup for evaluating several multi-omics models trained on pre-processed, harmonised and artificially horizontally federated data using Flower. Their study identifies a general but not substantial reduction in FL performance compared to centrally trained models, which increases with the number of clients and is variably affected by client heterogeneity.
@@ 4.5 Medical imaging
Medical imaging studies the human body’s interior to diagnose abnormalities in its anatomy and physiology from digital images such as those obtained by radiography, magnetic resonance and ultrasound devices (98). It is the most common application of FL in the medical literature (16). As a result, protocols for image segmentation and diagnostic prediction are well documented. Notable case studies target breast cancer (10), melanomas (83), cardiovascular disease (84), COVID-19 (85, 11).
Machine learning applications that use medical imaging data typically face challenges, including incomplete or inaccurate labelling and the normalisation of images from different scanners and different protocols. Bdair et al. (80) explored a federated labelling scheme in which clients produced ground-truth labels for skin lesions in a privacy-preserving manner, improving classification accuracy. Yan et al. (81) also proposed an efficient scheme to use data sets mainly comprising unlabelled images, focusing on chest X-rays. Furthermore, Jiang et al. (82) apply FL to learn a harmonised feature set from heterogeneous medical images, improving both the classification and segmentation of histology and MRI scans.
@@ 4.6 Ready-to-use FL tools for bioinformatics
The need for user-friendly FL implementations of common bioinformatics workflows has driven the creation of secure collaborative analysis tools (87, 88, 99). Two notable examples are sfkit and FeatureCloud.
@@
The sfkit framework (19) facilitates federated genomic analyses by implementing GWAS, principal component analysis (PCA), genetic relatedness and a modular architecture to complement them as needed. It provides a web interface featuring a project bulletin board, chat functions, study parameter configurations and results sharing. State-of-the-art cryptographic tools for privacy preservation based on SMPC and HE ensure data protection (100).
@@
FeatureCloud (86) is an integrated solution that enables end users without programming experience to build custom workflows. It provides modules to run on the clients and servers in the consortium. Unlike sfkit, FeatureCloud allows users to publish applications in its app store, including regression models, random forests and neural networks. Developers must also document how privacy guarantees are implemented in their apps.
5 Practical insights on federation
@@
This section offers practical insights to help readers interested in building a federated and secure analogue of an existing bioinformatics algorithm. We focus on horizontal FL with the centralised topology from Figure 2 (left). Consider different clients, each possessing a local data set , where . Each data set contains samples, denoted as , where represents the sample index, and represents the features for each sample. We denote a row (column) of the matrix as (). This describes a distributed data set of observations:
@@
The following sections assume that an FL consortium has been established, the necessary infrastructure is operational, and an appropriate FL framework has been selected and installed. It is also assumed that a secure aggregation protocol has been chosen, such as those described in Section 3.6 and Figure 4. The choice of a specific secure aggregation protocol may depend on several factors, including technology and infrastructure (e.g., the availability of a particular FL topology that drives the choice), as well as privacy risks and scalability concerns, as discussed in Section 3.6. In the following sections, we provide a general overview of sum-based mathematical operations built upon a secure aggregation protocol, as well as operations involving federated averaging (FedAvg; 6).
@@
Coding examples using Flower (31) are available in our GitHub repository (https://github.com/IDSIA/FL-Bioinformatics). We chose Flower because it has a shallow learning curve for new FL users and provides a good balance between simplicity and flexibility when implementing custom algorithms. Riedel et al. (28) also identified Flower as a promising framework because it has a large, active, and growing community of developers and scientists, as well as extensive tutorials and documentation. In our examples, secure summation is performed using the secure aggregation protocol SecAgg+ (101). This protocol combines encryption with SMPC, using a multiparty approach in which each client interacts with only a subset of the others. It is particularly suitable for several FL contexts, as it is robust to client dropout and highly scalable. In particular, a relevant aspect of the bioinformatics domain is that it scales linearly with the size of the vectors to be aggregated (102).
@@ 5.1 Sum-based computations
Let be real numbers stored by individual clients. We define the secure sum of these numbers, performed through the selected secure aggregation protocol, as . We can build on this simple, secure sum to construct a wide range of operations. However, note that as the complexity of operations increases, the amount of information revealed to the server may also increase. Sum-based operations include: